How much water should be added to 30 l of a 40% acid solution to reduce it to a 30% solution ?

Answer by richwmiller(17219)

from view of acid
30*0.4=0.3*(30+x)
12=9+0.3*x
12-9=0.3*x
3=0.3*x
x=10 liters of water
check
30*0.4=0.3*(30+x)
30*0.4=0.3*(40)
12=12
ok
or from view of water
1*y+0.6*30=0.7(30+y)
1*y+18=21+0.7y
1*y-0.7y=21-18
0.3y=3
y=3/0.3
y=10 liters of water
acid view x 10= water view y 10
match 10=10